# Asymmetric Errors

### v1

@@ 2.1.1 Flipped distributions
It can happen that an OPAT analysis results in both deviations having the same sign, as Figure 4.
Figure 4: The top row shows an OPAT analysis where the deviations are both positive, and the dimidiated and distorted fits. The second plot shows the dimidiated distribution: the total (black) is the sum of the two half-Gaussians (green and blue). The third plot also shows this pdf, together with the pdf obtained from the same 3 points using the distorted model, in green, and that from the equivalent pdf in red. The bottom row shows the equivalent for two negative deviations.
@@
Another scenario in which this formalism might be applied can occur with a discrete nuisance parameter; for example one might need to consider both the (favoured) normal and the (disfavoured but possible) inverted neutrino mass hierarchy. Suppose a result $R_{1}$ is obtained from the basic model or assumption, whereas a less-favoured alternative gives $R_{2}$. If, say $R_{2}>R_{1}$, one might wish to quote this as ${R_{1}}^{+R_{2}-R_{1}}_{-0}$. We are not recommending this, but if someone chooses to do so we can accommodate it as a flipped pdf where ${\sigma^{-}}$ and ${\sigma^{+}}$ are equal and given by $|R_{1}-R_{2}|$.
@@ 2.1.2 Comparison of the models
Name Description Range of Handles Notes
asymmetry flipping?
Dimidiated OPAT with 2 straight lines Limited but Special Discontinuity
generous case
Distorted OPAT with a parabola Any Yes Limited range
Railway OPAT with parabola Any Yes Arbitrary
morphing to straight lines smoothing
Double cubic OPAT with two cubics Any Yes Arbitrary
morphing to straight lines smoothing
Symmetric Beta OPAT with polynomial Any Yes Two arbitrary
morphing to straight lines tuning parameters
Quantile Variable Gaussian with $\sigma$ linear Any No Messy
Width function of cumulant numerically
Fechner Two half Gaussians Limited No Little
of same height motivation
Edgeworth Edgeworth expansion Very limited No Goes negative
Skew Normal Azzalini’s form Limited No
Maximum entropy Johnson $S_{U}$ Limited No kurtosis from
Johnson maximum entropy
Log-normal Takes log of Gaussian Any No Limited range
distributed variable
Table 1: Functions that can model asymmetric pdfs
@@
Figure 5: Pdfs for the different models. Panel A is for ${5.0}^{+1.1}_{-0.9}$, B is for moments: $\mu=5.0,V=2.0,\gamma=-1.0$, C is for ${5.0}^{+1.3}_{-0.7}$ and panel D for $\mu=5,V=2.0,\gamma=-3.0$
@@
Other examples can be tried, and show the same behaviour: for moderate skewness the functions are well behaved and results are similar; for large skewness they need to be handled carefully. The dimidiated Gaussian can accommodate arbitrary large values for ${{\sigma^{+}}-{\sigma^{-}}\over{\sigma^{+}}+{\sigma^{-}}}$, but not very large values of the normalised skewness $\gamma/V^{3/2}$.
@@ 2.2 Combination of pdf errors
The combination of two (or more) errors requires the convolution of the relevant pdfs: given $p_{x}(x)$ and $p_{y}(y)$ with $u=u(x,y)$ we require $p_{u}(u)$. In the linear approximation we assume, without loss of generality, that the values are shifted to have a nominal value of 0, $u(x,y)=u(x_{0},y_{0})+\tilde{u}(x-x_{0},y-y_{0})$ and then scaled to absorb the differentials, $x\to x_{0}+\left({\partial u\over\partial x}\right)\tilde{x}$ and $y\to y_{0}+\left({\partial u\over\partial y}\right)\tilde{y}$. Then (dropping the tildes) we have $u=x+y$ and the pdf for $u$ is
[…]
In some cases this can be done algebraically, in others it can be done numerically.
@@
This presents a problem, in that, although the convolution of two Gaussians gives another Gaussian, this is not in general true for other functions. The convolution of two pdfs, both belonging to one of the parametrisations in Table 1, does not give a function described by that parameterisation. See Appendix 22 for a detailed discussion using the dimidiated Gaussian.
@@
     To proceed, we note that the moments $\mu,V,\gamma$, add under convolution. To provide a consistent procedure for ‘adding errors’ we can evaluate the individual moments using Equations (A.1), (27), or their equivalents for other models, sum them to obtain the totals, and then use Equations (21), (28), or their equivalents to find the model parameters and, if desired, the quantile parameters, that give a pdf that has these moments.
$\sigma_{x}^{-}=-1.0$, $\sigma_{x}^{+}=1.0$, $\sigma_{y}^{-}=-0.8$, $\sigma_{y}^{+}=1.2$ Dimidiated Gaussian Distorted Gaussian Double Cubic Gaussian Edgeworth Expansion Fechner Distribution $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.32 1.52 0.080 1.33 1.54 0.098 1.32 1.52 0.080 1.31 1.52 0.095 1.30 1.52 0.092 Johnson System Log Normal QVW Gaussian Railway Gaussian Skew Normal $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.24 1.68 -0.006 1.33 1.59 0.075 1.32 1.52 0.083 1.34 1.53 0.098 1.33 1.49 0.101 Symmetric Beta Gaussian (1,10) Symmetric Beta Gaussian (1,15) Symmetric Beta Gaussian (1,20) Symmetric Beta Gaussian (1,25) Symmetric Beta Gaussian (1,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.32 1.52 0.080 1.33 1.53 0.087 1.33 1.53 0.092 1.33 1.53 0.094 1.33 1.53 0.096 Symmetric Beta Gaussian (2,10) Symmetric Beta Gaussian (2,15) Symmetric Beta Gaussian (2,20) Symmetric Beta Gaussian (2,25) Symmetric Beta Gaussian (2,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.32 1.52 0.079 1.32 1.52 0.083 1.33 1.53 0.088 1.33 1.53 0.091 1.33 1.53 0.093 Symmetric Beta Gaussian (2,10) Symmetric Beta Gaussian (2,15) Symmetric Beta Gaussian (2,20) Symmetric Beta Gaussian (2,25) Symmetric Beta Gaussian (2,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.32 1.52 0.078 1.32 1.52 0.081 1.32 1.52 0.085 1.33 1.53 0.089 1.33 1.53 0.091 Symmetric Beta Gaussian (4,10) Symmetric Beta Gaussian (4,15) Symmetric Beta Gaussian (4,20) Symmetric Beta Gaussian (4,25) Symmetric Beta Gaussian (4,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.32 1.52 0.078 1.32 1.52 0.080 1.32 1.52 0.084 1.33 1.53 0.087 1.33 1.53 0.090 $\sigma_{x}^{-}=-0.8$, $\sigma_{x}^{+}=1.2$, $\sigma_{y}^{-}=-0.8$, $\sigma_{y}^{+}=1.2$ Dimidiated Gaussian Distorted Gaussian Double Cubic Gaussian Edgeworth Expansion Fechner Distribution $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.22 1.62 0.160 1.25 1.64 0.203 1.23 1.62 0.160 1.21 1.62 0.195 1.19 1.62 0.186 Johnson System Log Normal QVW Gaussian Railway Gaussian Skew Normal $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.28 1.83 0.156 1.26 1.70 0.188 1.23 1.62 0.167 1.25 1.64 0.199 1.22 1.60 0.190 Symmetric Beta Gaussian (1,10) Symmetric Beta Gaussian (1,15) Symmetric Beta Gaussian (1,20) Symmetric Beta Gaussian (1,25) Symmetric Beta Gaussian (1,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.23 1.62 0.162 1.23 1.63 0.176 1.24 1.63 0.186 1.24 1.63 0.192 1.24 1.64 0.195 Symmetric Beta Gaussian (2,10) Symmetric Beta Gaussian (2,15) Symmetric Beta Gaussian (2,20) Symmetric Beta Gaussian (2,25) Symmetric Beta Gaussian (2,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.23 1.62 0.159 1.23 1.62 0.168 1.23 1.63 0.178 1.24 1.63 0.185 1.24 1.63 0.190 Symmetric Beta Gaussian (2,10) Symmetric Beta Gaussian (2,15) Symmetric Beta Gaussian (2,20) Symmetric Beta Gaussian (2,25) Symmetric Beta Gaussian (2,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.22 1.62 0.158 1.23 1.62 0.164 1.23 1.63 0.173 1.23 1.63 0.180 1.24 1.63 0.185 Symmetric Beta Gaussian (4,10) Symmetric Beta Gaussian (4,15) Symmetric Beta Gaussian (4,20) Symmetric Beta Gaussian (4,25) Symmetric Beta Gaussian (4,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.22 1.62 0.157 1.23 1.62 0.162 1.23 1.62 0.169 1.23 1.63 0.176 1.23 1.63 0.182 $\sigma_{x}^{-}=-0.5$, $\sigma_{x}^{+}=1.5$, $\sigma_{y}^{-}=-0.8$, $\sigma_{y}^{+}=1.2$ Dimidiated Gaussian Distorted Gaussian Double Cubic Gaussian Edgeworth Expansion Fechner Distribution $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.09 . 1.78 . 0.284 1.17 . 1.88 . 0.349 1.11 . 1.79 . 0.285 Johnson System Log Normal QVW Gaussian Railway Gaussian Skew Normal $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 0.73 . 1.93 . 0.115 1.12 . 1.80 . 0.298 1.18 . 1.86 . 0.325 Symmetric Beta Gaussian (1,10) Symmetric Beta Gaussian (1,15) Symmetric Beta Gaussian (1,20) Symmetric Beta Gaussian (1,25) Symmetric Beta Gaussian (1,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.12 . 1.80 . 0.288 1.14 . 1.82 . 0.313 1.15 . 1.83 . 0.330 1.16 . 1.85 . 0.339 1.17 . 1.86 . 0.343 Symmetric Beta Gaussian (2,10) Symmetric Beta Gaussian (2,15) Symmetric Beta Gaussian (2,20) Symmetric Beta Gaussian (2,25) Symmetric Beta Gaussian (2,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.11 1.79 0.283 1.13 1.81 0.300 1.14 1.82 0.317 1.15 1.83 0.328 1.16 1.84 0.335 Symmetric Beta Gaussian (2,10) Symmetric Beta Gaussian (2,15) Symmetric Beta Gaussian (2,20) Symmetric Beta Gaussian (2,25) Symmetric Beta Gaussian (2,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.11 1.79 0.281 1.12 1.80 0.293 1.13 1.81 0.308 1.14 1.82 0.320 1.15 1.83 0.328 Symmetric Beta Gaussian (4,10) Symmetric Beta Gaussian (4,15) Symmetric Beta Gaussian (4,20) Symmetric Beta Gaussian (4,25) Symmetric Beta Gaussian (4,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.11 1.79 0.280 1.12 1.80 0.288 1.13 1.81 0.302 1.14 1.82 0.314 1.15 1.83 . 0.323 $\sigma_{x}^{-}=-0.5$, $\sigma_{x}^{+}=1.5$, $\sigma_{y}^{-}=-0.5$, $\sigma_{y}^{+}=1.5$ Dimidiated Gaussian Distorted Gaussian Double Cubic Gaussian Edgeworth Expansion Fechner Distribution $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 0.97 1.93 0.413 1.12 2.07 0.534 1.00 1.95 0.419 Johnson System Log Normal QVW Gaussian Railway Gaussian Skew Normal $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 0.91 2.42 0.351 1.03 1.96 0.440 1.13 2.04 0.484 Symmetric Beta Gaussian (1,10) Symmetric Beta Gaussian (1,15) Symmetric Beta Gaussian (1,20) Symmetric Beta Gaussian (1,25) Symmetric Beta Gaussian (1,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.01 1.95 0.423 1.05 1.98 0.466 1.08 2.00 0.495 1.10 2.02 0.511 1.11 2.03 0.519 Symmetric Beta Gaussian (2,10) Symmetric Beta Gaussian (2,15) Symmetric Beta Gaussian (2,20) Symmetric Beta Gaussian (2,25) Symmetric Beta Gaussian (2,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.00 1.95 0.415 1.03 1.96 0.443 1.06 1.98 0.472 1.08 2.00 0.492 1.09 2.01 0.504 Symmetric Beta Gaussian (2,10) Symmetric Beta Gaussian (2,15) Symmetric Beta Gaussian (2,20) Symmetric Beta Gaussian (2,25) Symmetric Beta Gaussian (2,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 1.00 1.95 0.411 1.02 1.96 0.431 1.04 1.97 0.457 1.06 1.99 0.478 1.08 2.00 0.492 Symmetric Beta Gaussian (4,10) Symmetric Beta Gaussian (4,15) Symmetric Beta Gaussian (4,20) Symmetric Beta Gaussian (4,25) Symmetric Beta Gaussian (4,30) $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ 0.99 1.94 0.409 1.01 1.95 0.424 1.03 1.97 0.446 1.05 1.98 0.466 1.07 1.99 0.482
@@
Table 2: Combining errors from various pdfs using various models. - ARB: $\Delta$ is combined median - 00
@@
Table 2 illustrates the results of combining various errors using different models. The skew normal fails to accommodate the examples with very high skewness. The behaviour is much as one would expect. The combined errors increase in all cases, but the asymmetry falls. This is brought out in the bottom entry, where two very asymmetric errors (${\sigma^{+}}=3{\sigma^{-}}$) combine to give an error with only ${\sigma^{+}}\approx 2{\sigma^{-}}$. For moderate asymmetries, as in the first two rows, there is reasonable agreement between the models, to two significant figures, but not three, which gives an estimate of how far the precision of such calculations can be trusted. For large asymmetries, as in the bottom two rows, the agreement is lost, showing that in such cases the accuracy should not be considered as being definite. After all, any method will break down somewhere.
@@
An important point to note is that in all cases the nominal value shifts: this is the number shown as $\Delta$. If two asymmetric pdfs are convolved, then the median of the result is not the sum of the individual medians. A procedure for combining errors using asymmetric distributions must, to be consistent, include a shift in the central value. If practitioners are not prepared to make such an adjustment then they should perhaps reconsider the use of asymmetric errors and revert to the usual symmetric form, averaging their ${\sigma^{+}}$ and ${\sigma^{-}}$ values.

### v2

@@ 2.1.1 Flipped distributions
It can happen that an OPAT analysis results in both deviations having the same sign, as illustrated in Figure 5.
Figure 5: The top row shows an OPAT analysis where the deviations are both positive, though of differing magnitudeand the dimidiated and distorted fits. The second plot shows the dimidiated distribution: the total (black) is the sum of the two half-Gaussians (green and blue). The third plot also shows this pdf, together with the pdf obtained from the same 3 points using the distorted model, in green, and that from the equivalent pdf in red. The bottom row shows the equivalent for two negative deviations.
@@
Another scenario in which this formalism might be applied can occur with a discrete nuisance parameter; for example one might need to consider both the (favoured) normal and the (disfavoured but possible) inverted neutrino mass hierarchy. Suppose a result $R_{1}$ is obtained from the basic model or assumption, whereas a less-favoured alternative gives $R_{2}$. If, say $R_{2}>R_{1}$, one might wish to quote this as ${R_{1}}^{+R_{2}-R_{1}}_{-0}$. We are not recommending this, but if someone chooses to do so we can accommodate it as a flipped pdf where ${\sigma^{-}}$ and ${\sigma^{+}}$ are equal (or opposite: the signs require careful handling in the software) and given by $|R_{1}-R_{2}|$.
@@ 2.1.2 Comparison of the models
The Asymmetry $A$ is here defined as ${{\sigma^{+}}-{\sigma^{-}}\over{\sigma^{+}}+{\sigma^{-}}}$. Many models can handle arbitrary values of ${\sigma^{+}}$ and ${\sigma^{-}}$ and thus arbitrary $A$. Others, such as the Fechner and the skew normal, for which the extreme case is a half-Gaussian, cannot model the largest $|A|$. The dimidiated Gaussian can handle arbitrary $A$ but not arbitrary $\gamma/V^{3/2}$.
Name Description Range of $A$ Handles flipping? Notes
Dimidiated Gaussian OPAT with 2 straight lines ±$\pm 1$ Special case Discontinuity which can give problems when fitting
Distorted Gaussian OPAT with a parabola ±$\pm 0.57382$ Yes Limited range
Railway Gaussian OPAT with parabola morphing to straight lines ±$\pm 0.60467$ Yes Arbitrary smoothing
Double cubic Gaussian OPAT with two cubics morphing to straight lines ±$\pm 0.74538$ Yes Arbitrary smoothing
Symmetric beta Gaussian OPAT with polynomial morphing to straight lines ±$\pm 1$ Yes Two arbitrary tuning parameters
QVW Gaussian Gaussian with $\sigma$ linear function of cumulant ±$\pm 0.68269$ No Messy numerically
Fechner distribution Two half Gaussians of same height ±$\pm 0.21564$ No Little motivation
Edgeworth expansion Edgeworth expansion Very limited No Goes negative
Skew Normal Azzalini’s form ±$\pm 0.21564$ No
Maximum entropy Johnson Johnson $S_{U}$$\pm 0.34719$ No Kurtosis from maximum entropy
Log normal Takes exp of Gaussian distributed variable ±$\pm 1$ No Limited range
Table 2: Functions that can model asymmetric pdfs.
@@
Figure 6: Pdfs for the different models. Panel A is for ${5.0}^{+1.1}_{-0.9}$, B is for moments: $\mu=5.0,V=2.0,\gamma=-1.0$, C is for ${5.0}^{+1.3}_{-0.7}$ and panel D for $\mu=5,V=2.0,\gamma=-3.0$.
@@
Other examples can be tried, and show the same behaviour: for moderate skewness the functions are well behaved and results are similar; for large skewness they need to be handled carefully.
@@ 2.2 Combination of pdf errors
The combination of two (or more) errors requires the convolution of the relevant pdfs: given $p_{x}(x)$ and $p_{y}(y)$ with $u=u(x,y)$ we require $p_{u}(u)$. In the linear approximation we assume, without loss of generality, that the values are shifted to have a nominal value of 0, $u(x,y)=u(x_{0},y_{0})+\tilde{u}(x-x_{0},y-y_{0})$ and then scaled to absorb the differentials, $x\to x_{0}+\left({\partial u\over\partial x}\right)\tilde{x}$ and $y\to y_{0}+\left({\partial u\over\partial y}\right)\tilde{y}$. Then (dropping the tildes) we have $u=x+y$ and the pdf for $u$ is the convolution of the two , as given in Equation 3.In some cases this can be done algebraically, in others it can be done numerically.
@@
This presents a problem, in that, although the convolution of two Gaussians gives another Gaussian, this is not in general true for other functions. The convolution of two pdfs, both belonging to one of the parametrisations in Table 2, does not give a function described by that parameterisation. See Appendix A.1.1 for a detailed discussion using the dimidiated Gaussian.
@@
     To proceed, we note that the moments $\mu,V,\gamma$, add under convolution. To provide a consistent procedure for ‘adding errors’ we can evaluate the individual moments using Equations (17), (25), or their equivalents for other models, sum them to obtain the totals, and then use Equations (19), (26), or their equivalents to find the model parameters and, if desired, the quantile parameters, that give a pdf that has these moments.
Errors
$\sigma_{x}^{-}=1.0$, $\sigma_{x}^{+}=1.0$ $\sigma_{x}^{-}=0.8$, $\sigma_{x}^{+}=1.2$ $\sigma_{x}^{-}=0.5$, $\sigma_{x}^{+}=1.5$ $\sigma_{x}^{-}=0.5$, $\sigma_{x}^{+}=1.5$
$\sigma_{y}^{-}=0.8$, $\sigma_{y}^{+}=1.2$ $\sigma_{y}^{-}=0.8$, $\sigma_{y}^{+}=1.2$ $\sigma_{y}^{-}=0.8$, $\sigma_{y}^{+}=1.2$ $\sigma_{y}^{-}=0.5$, $\sigma_{y}^{+}=1.5$
Models $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$ $\sigma^{-}$ $\sigma^{+}$ $\Delta$
Dimidiated Gaussian 1.32 1.52   0.080 1.22 1.62 0.160 1.09 1.78 0.284 0.97 1.93 0.413
Distorted Gaussian 1.33 1.54   0.098 1.25 1.64 0.203 1.17 1.88 0.349 1.12 2.07 0.534
Railway Gaussian 1.34 1.53   0.098 1.25 1.64 0.199 1.18 1.86 0.325 1.13 2.04 0.484
Double cubic Gaussian 1.32 1.52   0.080 1.23 1.62 0.160 1.11 1.79 0.285 1.00 1.95 0.419
Symmetric beta Gaussian (1,10) 1.32 1.52   0.080 1.23 1.62 0.162 1.12 1.80 0.288 1.01 1.95 0.423
Symmetric beta Gaussian (1,30) 1.33 1.53   0.096 1.24 1.64 0.195 1.17 1.86 0.343 1.11 2.03 0.519
Symmetric beta Gaussian (4,10) 1.32 1.52   0.078 1.22 1.62 0.157 1.11 1.79 0.280 0.99 1.94 0.409
Symmetric beta Gaussian (4,30) 1.33 1.53   0.090 1.23 1.63 0.182 1.15 1.83 0.323 1.07 1.99 0.482
QVW Gaussian 1.32 1.52   0.083 1.23 1.62 0.167 1.12 1.80 0.298 1.03 1.96 0.440
Fechner distribution 1.30 1.52   0.092 1.19 1.62 0.186
Edgeworth expansion 1.31 1.52   0.095 1.21 1.62 0.195
Skew normal 1.33 1.49   0.101 1.22 1.60 0.190
Johnson system 1.24 1.68 $-0.006$ 1.28 1.83 0.156
Log normal 1.33 1.59   0.075 1.26 1.70 0.188 0.73 1.93 0.115 0.91 2.42 0.351
@@
Table 3: Combining errors from various pdfs using various models. $\Delta$ is the combined median.
@@
Table 3 illustrates the results of combining various typical errors, using different models. The behaviour is much as one would expect. The combined errors increase in all cases, but the asymmetry falls. This is brought out in the rightmost entry, where two very asymmetric errors (${\sigma^{+}}=3{\sigma^{-}}$) combine to give an error with only ${\sigma^{+}}\approx 2{\sigma^{-}}$. For moderate asymmetries, as in the first two columns, there is reasonable agreement between the models, to two significant figures, but not three, which gives an estimate of how far the precision of such calculations can be trusted. For large asymmetries, as in the third and fourth columns, the agreement is lost, showing that in such cases the accuracy should not be considered as being definite. (Several models fail to accommodate the examples with very high asymmetries, so the entries are blank). After all, any method will break down somewhere.
@@
An important point to note is that in all cases the nominal value shifts: this is the number shown as $\Delta$. If two asymmetric pdfs are convolved, then the median of the result is not the sum of the individual medians. A procedure for combining errors using asymmetric distributions must, to be consistent, include a shift in the central value. If practitioners are not prepared to make such an adjustment then they should perhaps reconsider the use of asymmetric errors and revert to the usual symmetric form, averaging their ${\sigma^{+}}$ and ${\sigma^{-}}$ values. The Values of ${\sigma^{+}}$ and ${\sigma^{-}}$ in the table, and returned by the software, give, within the framework of the model concerned, the 68% central region around the nominal point: when the $\Delta$ shift is applied they should be adjusted accordingly.
2.3 Why the ‘usual procedure’ is wrong
The flaw in the logic is as follows: when two random variables, described (for the sake of the example) by dimidiated Gaussians, are combined, there is a 25% chance that they will both fluctuate upwards. That is indeed described, within the framework of this model, by a 2-D Gaussian for which the standard deviation of the sum is the sum in quadrature of the two positive $\sigma$. Likewise there is a 25% chance that they will both go negative, and that is described by $\sqrt{{\sigma^{-}_{1}}^{2}+{\sigma^{-}_{2}}^{2}}$. But there is a 50% chance that one will go up and the other will go down. This is neglected by the ‘usual procedure’, and it is this filling in of central values that lets the Central Limit Theorem do its work.
Examples of how this can give wrong results are in Section 4.1.2.